OpenStudy (anonymous):

please help! i don't understand this. If sin0=2/3, find the values of the other five trigonometric functions of 0. (the 0's have a slashed line through it)

OpenStudy (anonymous):

i think in question it should be sin x.....please check it

OpenStudy (anonymous):

no it says sin0, but the 0 has a dash through if. do you know what i mean by that?

OpenStudy (anonymous):

it looks like the one on the top right corner

OpenStudy (anonymous):

@erin512

OpenStudy (anonymous):

YA ITS right now. the symbol is pronounced theta

OpenStudy (anonymous):

oh okay sorry

OpenStudy (anonymous):

\[\sin \theta =2/3\] \[\sin \theta=P/H\] where P stands for perpendicular and H stands for hypotenuse B denotes the base By pythagoras theorem... H^2=B^2+P^2 \[H=\sqrt{(P)^{2}+(B)^{2}}\] FROM THIS.. \[B=\sqrt{(H)^{2}-(P)^{2}}\] \[B=\sqrt{9-4}\] \[B=\sqrt{5}\]

OpenStudy (anonymous):

\[\cos \theta=B/H\] \[\cos \theta=\sqrt{5}/3\]

OpenStudy (anonymous):

\[\tan \theta=P/B\] \[\tan \theta=2/\sqrt{5}\]

OpenStudy (anonymous):

\[\sec \theta=1/(\cos \theta)\] \[\cot \theta=1/(\tan \theta)\] \[\csc \theta=1/(\sin \theta)\]

OpenStudy (anonymous):

i hope u can find them now......

OpenStudy (anonymous):

what am i trying to find now? thanks for all the help. i'm sorry, i'm so lost in this unit

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

put the values of \[\cos \theta,\tan \theta,\sin \theta \] in THE ABOVE THREE expressions to find the values of these three angles respectively...

OpenStudy (anonymous):

not these three angles......it should be other three angles..... ie \[\sec \theta ,\cot \theta,\csc \theta \]

OpenStudy (anonymous):

cosθ,tanθ,sinθ = 0.75, 0.89, 0.67. is this correct?

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

i'm not sure what the final answer would be

OpenStudy (anonymous):

oh are those the first 3 values its looking for?